Derive expressions for potential energy and kinetic energy of an oscillating spring-
mass system
1
Expert's answer
2019-04-01T10:39:49-0400

The equation of harmonic motion:


"x = A\sin(\omega t + \phi_0)" , where "\omega = \sqrt{\frac{k}{m}}" in case oscillating spring-mass system.

"\phi_0" - initial oscillation phase (actually if it is, for example, "-\frac{\pi}{4}" we can rewrite it with cosine)

"\dot x = A \cdot \cos(\omega t + \phi_0) \cdot \omega"


"E_k = \frac{mv^2}{2} = \frac{m\dot x^2}{2} = \frac{m(A\omega \cos(\omega t + \phi_0))^2}{2} = \frac{m\omega^2(A \cos(\omega t + \phi_0))^2}{2} =
\frac{m\frac{k}{m}(A \cos(\omega t + \phi_0))^2}{2} \Rightarrow"



"E_k = \frac{k(A\cos(\omega t + \phi_0))^2}{2} = \frac{k(A\cos(\sqrt{\frac{k}{m}} t + \phi_0))^2}{2}"

"E_p = \frac{kx^2}{2} \Rightarrow"



"E_p = \frac{k(A\sin(\omega t + \phi_0))^2}{2} = \frac{k(A\sin(\sqrt{\frac{k}{m}} t + \phi_0))^2}{2}"

As we can see expressions differ only in the initial phase (we can change the sine to cosine by changing the initial phase). with Pythagorean trigonometric identity:

"E_{mechanical} = E_p + E_k = \frac{kA^2}{2} = const"